OpenStudy (anonymous):

Which is the solution of the equation below, rounded to the nearest tenth? 2.1x – 4.09 = 6.98x + 1.73

OpenStudy (anonymous):

i really need help please

OpenStudy (anonymous):

Convert the decimal numbers to their equivalent fractions and then solve the resulting equation for x. The answer for x will be in exact symbolic form.\[\frac{21}{10}x -\frac{409}{100} \text{=} \frac{698}{100}x +\frac{173}{100}\]\[x=-\frac{291}{244} \]x to 40 digits.\[-1.192622950819672131147540983606557377049 \]

OpenStudy (anonymous):

Xavier's dining room table was unstable because one leg was shorter than the others. He placed a 3/8 inch thick piece of wood under the short leg, and it became even with the other legs, which are 19 1/4 inches long. How many inches long was the short leg before Xavier added the wood? 19 5/8 inches 18 3/4 inches 19/ 7/8 inches 18 7/8 inches

OpenStudy (anonymous):

so it would be -1.1

OpenStudy (anonymous):

@robtobey plz help Xavier's dining room table was unstable because one leg was shorter than the others. He placed a 3/8 inch thick piece of wood under the short leg, and it became even with the other legs, which are 19 1/4 inches long. How many inches long was the short leg before Xavier added the wood? 19 5/8 inches 18 3/4 inches 19/ 7/8 inches 18 7/8 inches

OpenStudy (anonymous):

happy face get outta here go ask ur question somewhere eles i need help

OpenStudy (anonymous):

@budsmoker101 -1.19 when rounded to the nearest 0.01 .

OpenStudy (anonymous):

o.4 0or 0.5 i will give u a medal

OpenStudy (anonymous):

@budsmoker101 x = -1.19262 when the equation is solved using the decimal numbers. You have to decide on the precision in the answer. x = -1.2 or x = -1.19 or x = -1.193 and so on.